# Refer to Anything with Vision-Language Prompts

### v1

@@ 1 Introduction
Referring expression segmentation (RES) [69, 34, 18] enables language-based object segmentation by relating a text prompt with a segmentation mask for the referred target, and can be generalized (i.e., GRES [34]) for multiple targets. However, real-world applications, such as autonomous driving [52, 9], robotics [14, 37, 64], augmented reality [25], and image editing [53, 41, 15], require additional flexibility in the prompt. As exemplified in Figure 1, a user may want to specify a segmentation instruction that involves a relationship/comparison/interaction with a reference visual entity (see red regions in the middle column). Expressing the reference via visual and text prompts together is usually preferred over text-only descriptions in such cases, because language may not be able to concisely and accurately locate the reference entity or describe its intricate characteristics in a complex image.
In this work, we introduce a novel task of omnimodal referring expression segmentation (ORES) with vision-language prompts (Figure 1). For a given image, the system generates a group of relevant masks that satisfy a user-specified instruction, which can be a) a text-only prompt describing a property (e.g., category, attribute, position, or their combination) of the targets, or b) a vision-enriched prompt that provides masks of reference entities and expresses a complex property involving the reference entities.
@@ 2 Related Work
Large multimodal models (LMMs) extend large language models (LLMs) [11, 47, 58] with vision-language capabilities via visual instruction tuning [35, 75, 10]. Early LMMs are mainly based on CLIP [48] patch-level visual features and show weaknesses in object-level comprehension and reasoning [57, 56, 30, 55]. LMMs can be equipped with grounding capabilities for generating bounding boxes [43, 4, 60, 44, 68, 28] or segmentation masks [27, 49, 73, 50] via training on converted datasets with semantic-pixel alignment. Unlike prior LMMs, our RAS is not trained for text generation, because text responses are unnecessary in the task and applications we consider (Figure 1), and high-quality mask groups are prioritized.
@@ 4.2 MaskGroups-HQ: Human-Annotated Data for Finetuning and Evaluation
Although MaskGroups-2M is large enough for instruction tuning, its mask groups in pre-defined templates cannot cover all possible criteria that human users may use for grouping, and may introduce noises due to inaccurate annotations in the source data. Therefore, to further improve and evaluate the generalizability of RAS, we manually annotate a high-quality dataset MaskGroups-HQ (visualized in Figure 3). We start from EntitySeg [46], an image segmentation dataset containing high-resolution images and mask annotations for category-agnostic visual entities within images. Then, human annotators inspect the images and masks, and annotate several mask groups by proposing a reasonable vision-language prompt and labeling the IDs of masks that should be included. Our quality check ensures that the proposed mask groups are agreed upon by different users. In total, 100,299 mask groups are annotated. Finally, we split MaskGroups-HQ into 96,697 samples over 18,368 images for finetuning, and 3,599 samples over 661 images for evaluation. 28% of the samples include reference mask(s) in their prompts.
@@ 5.1 Omnimodal RES
Our new ORES task poses new challenges to existing GRES models, because it uses mask-based visual prompts to describe relationships with reference entities. In Table 3, we compare RAS (both before and after finetuning on MaskGroups-HQ) based on off-the-shelf SAM [24] for proposing candidate masks, with state-of-the-art GRES models [34, 74, 62]. Unlike all GRES baselines, RAS is able to accept reference masks and process all prompts in MaskGroups-HQ, demonstrating a stronger prompting flexibility. Meanwhile, for text-only prompts, our results show significantly improved generalizability. Further finetuning RAS on MaskGroups-HQ brings more performance gains, especially enhancing the ability for understanding reference masks. Quantitative results are shown in Figure C in the supplementary material.
@@ 5.3 Analysis and Ablation Study
Quality of candidate masks. RAS extends existing segmentation models with complex vision-language interactions. To validate that the segmentation models can propose candidate masks that sufficiently cover the referred targets, we analyze the proposed masks by selecting the ones with the hightest IoU with ground-truth masks and computing their cIoU on all three tasks. As shown in Table 6, even without finetuning the segmentation models on the RES tasks, the best candidates achieve significantly better cIoU than existing models, and the proposals indeed include most of the true targets (over 85 cIoU). By addressing the challenge of semantically understanding the high-quality candidate masks, RAS achieves the best final performance.
@@ Appendix D Additional Experiment Results
Converting visual prompts into language. In the main paper, we have discussed the limitations of existing GRES models [34, 62, 74]. They cannot take visual prompts that represent reference entities as inputs, and therefore cannot process all samples in the ORES task (Table 3). One may argue that visual prompts in ORES can be replaced by text prompts (e.g., “Locate all pillows on <mask-ref>” “Locate all pillows on the bed”, in the right column of Figure 1). However, when the scene is complex and involves multiple semantically similar objects, visual prompts can hardly be clearly and concisely “translated” into language. To investigate this discrepancy between visual prompts and text prompts, we manually convert <mask-ref> into language for 200 samples in MaskGroups-HQ, and test GRES models and our RAS on these samples. As shown in Table B, visual prompts are better perceived by RAS, indicating that such visual prompts are necessary to guide the model in accurately locating the target entities that are related to the reference entity. When provided with the same pure-text prompts, despite the increased complexity of the converted prompts, RAS still outperforms the existing GRES models.
@@ Appendix F Additional Qualitative Results
In Figure C, we provide additional visualized results of applying RAS and other GRES models in the ORES task. We observe that RAS (both before and after finetuned on MaskGroups-HQ) achieves better results on MaskGroups-HQ than all previous GRES models, which is consistent with our quantitative evaluation in Table 3 of the main paper.

### v2

# Refer to Any Segmentation Mask Group With Vision-Language Prompts

@@ 1 Introduction
Referring expression segmentation (RES) [69, 34, 18] enables language-based object segmentation by relating a text prompt with a segmentation mask for the referred target, and can be generalized (i.e., GRES [34]) for multiple targets. However, real-world applications, such as autonomous driving [52, 9], robotics [14, 37, 64], augmented reality [25], and image editing [53, 41, 15], require additional flexibility in the prompt. As exemplified in Figure LABEL:fig:teaser, a user may want to specify a segmentation instruction that involves a relationship/comparison/interaction with a reference visual entity (see red regions in the middle column). Expressing the reference via visual and text prompts together is usually preferred over text-only descriptions in such cases, because language may not be able to concisely and accurately locate the reference entity or describe its intricate characteristics in a complex image.
In this work, we introduce a novel task of omnimodal referring expression segmentation (ORES) with vision-language prompts (Figure LABEL:fig:teaser). For a given image, the system generates a group of relevant masks that satisfy a user-specified instruction, which can be a) a text-only prompt describing a property (e.g., category, attribute, position, or their combination) of the targets, or b) a vision-enriched prompt that provides masks of reference entities and expresses a complex property involving the reference entities.
@@ 2 Related Work
Large multimodal models (LMMs) extend large language models (LLMs) [11, 47, 58] with vision-language capabilities via visual instruction tuning [35, 75, 10]. Early LMMs are mainly based on CLIP [48] patch-level visual features and show weaknesses in object-level comprehension and reasoning [57, 56, 30, 55]. LMMs can be equipped with grounding capabilities for generating bounding boxes [43, 4, 60, 44, 68, 28] or segmentation masks [27, 49, 73, 50] via training on converted datasets with semantic-pixel alignment. Unlike prior LMMs, our RAS is not trained for text generation, because text responses are unnecessary in the task and applications we consider (Figure LABEL:fig:teaser), and high-quality mask groups are prioritized.
@@ 4.2 MaskGroups-HQ: Human-Annotated Data for Finetuning and Evaluation
Although MaskGroups-2M is large enough for instruction tuning, its mask groups in pre-defined templates cannot cover all possible criteria that human users may use for grouping, and may introduce noises due to inaccurate annotations in the source data. To further improve and evaluate the generalizability of RAS, we manually annotate a high-quality dataset MaskGroups-HQ (visualized in Figure 3). We start from EntitySeg [46], an image segmentation dataset containing high-resolution images and category-agnostic masks. Notably, the images are from various sources, not only MS-COCO. Then, human annotators inspect the images and masks, and annotate several mask groups by proposing a reasonable vision-language prompt and labeling the IDs of masks that should be included. Our quality check ensures that the proposed mask groups are agreed upon by different users. In total, 100,299 mask groups are annotated. Finally, we split MaskGroups-HQ into 96,697 samples over 18,368 images for finetuning, and 3,599 samples over 661 images for evaluation. 28% of the samples include reference mask(s) in their prompts.
@@ 5.1 Omnimodal RES
Our new ORES task poses new challenges to existing GRES models, because it uses mask-based visual prompts to describe relationships with reference entities. In Table 3, we compare RAS (both before and after finetuning on MaskGroups-HQ) based on off-the-shelf SAM [24] for proposing candidate masks, with state-of-the-art GRES models [34, 74, 62]. Unlike all GRES baselines, RAS is able to accept reference masks and process all prompts in MaskGroups-HQ, demonstrating a stronger prompting flexibility. Meanwhile, for text-only prompts, our results show significantly improved generalizability. Further finetuning RAS on MaskGroups-HQ brings more performance gains, especially enhancing the ability for understanding reference masks. Quantitative results are shown in Figure D in the supplementary material.
@@ 5.3 Analysis and Ablation Study
Quality of candidate masks. RAS extends existing segmentation models with complex vision-language interactions. To validate that the segmentation models, Co-DETR and SAM, can propose candidate masks that sufficiently cover the referred targets, we analyze the proposed masks by selecting the ones with the hightest IoU with ground-truth masks and computing their cIoU on all three tasks. As shown in Table 6, even without finetuning on RES tasks, the best candidates by Co-DETR and SAM achieve significantly better cIoU than existing RES models, and the proposals indeed include most of the true targets ( 85 cIoU). By addressing the challenge of understanding and selecting high-quality candidates, RAS achieves the best final results.
@@ 6 Conclusion
Acknowledgments.
This work was supported in part by NSF Grant 2106825 and NIFA Award 2020-67021-32799. This work used computational resources, including the NCSA Delta and DeltaAI supercomputers through allocations CIS230012, CIS230013, CIS240133, and CIS240428 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, as well as the TACC Frontera supercomputer, Amazon Web Services, and OpenAI API through the National Artificial Intelligence Research Resource (NAIRR) Pilot.
Appendix
In this appendix, we provide additional details on the implementation of our model RAS (Section A) and our datasets MaskGroups-2M (Section B) and MaskGroups-HQ (Section C). Furthermore, we include additional experiments (Section D), ablation study (Section E), and qualitative results (Section F).
@@ Appendix D Additional Experiment Results
Converting visual prompts into language. In the main paper, we have discussed the limitations of existing GRES models [34, 62, 74]: They cannot take visual prompts that represent reference entities as inputs, and therefore cannot process all samples in the ORES task (Table 3). One may argue that visual prompts in ORES can be replaced by text prompts (e.g., “Locate all pillows on <mask-ref>” “Locate all pillows on the bed”, Figure LABEL:fig:teaser). However, when the scene is complex and involves multiple semantically similar objects, visual prompts can hardly be clearly and concisely “translated” into language. To investigate this discrepancy between visual prompts and text prompts, we manually convert <mask-ref> into language for 200 samples in MaskGroups-HQ, and test GRES models and our RAS on these samples. As shown in Table B, visual prompts are better perceived by RAS, indicating that such visual prompts are necessary to guide the model in accurately locating the target entities that are related to the reference entity. When provided with the same pure-text prompts, despite the increased complexity of the converted prompts, RAS still outperforms the existing GRES models.
@@ Appendix F Additional Qualitative Results
Our RAS shows strong generalization beyond MS-COCO benchmarks, where prior works primarily focus. As shown in Figure C, our model outperforms GSVA on out-of-distribution (OOD) images. This is achieved by decoupling mask generation and selection, allowing RAS to leverage strong generalization capabilities of SAM.
In Figure D, we provide additional visualized results of applying RAS and other GRES models in the ORES task. RAS (both before and after finetuned on MaskGroups-HQ) achieves better results on MaskGroups-HQ than all previous GRES models, which is consistent with our quantitative evaluation in Table 3 of the main paper.